# CCC++: Optimized Color Classified Colorization with Segment Anything Model (SAM) Empowered Object Selective Color Harmonization

### v1

@@ III-B Problem Definition
The colorization problem is considered as predicting color channels from a given gray channel. The Lightness (L) channel of La*b* color space can be mapped into the gray channel (intensity) and vice versa[51]. Furthermore RGB can be mapped into LAB and vice versa. The task can be defined as follows in Eq. 1, 2, 3, and 4.
@@
where $X_{L}$ is the lightness channel, $X_{ab}$ is the predicted color channel, $X_{Lab}$ is the predicted color image, $Y_{ab}$ is the ground truth color channel, $f(.)$ is the mapping function achieved by deep learning, $Distance_{min}(.)$ is the objective function (can be any loss function) by which the optimizer makes the learning efficient, $\mathbb{R}$ is the total image component, $\mathbb{H}$ and $\mathbb{W}$ are the image dimension.
Bin Size = 14 Refer to caption Bin Size = 12 Refer to caption Bin Size = 10 Refer to caption Bin Size = 8 Refer to caption Bin Size = 6 Refer to caption Bin Size = 4
@@
Theoretically, the values of the a* and b* channel are continuous within [-128, 127]. So, the prediction is considered a regression problem. Therefore, the $Distance_{min}(.)$ of equation 2 naturally can be either L1 loss or L2 loss or Huber loss or Log-cosh loss or similar regression loss. Loss functions are shown below.
@@
where $N$ represents the total number of pixels in a batch and $\delta$ represents the acceptable threshold. Background colors such as clouds, soil, pavement, and walls cover the majority of the areas in the real-time images. The color components of the primary or focal objects are substantially smaller than the total color components. This results in an unbalanced distribution of features. Handling feature imbalance is crucial because the smaller subsets of features are the feature of interest for the learning task. The colorization problem’s inherent ambiguity and multimodality make the above loss functions vulnerable. If an object can take on a variety of a*b* values, the mean of the set is the most effective method to solve the loss shown in Eq. 5, 6, 7 and 8. The averaging error effect favors color values that are predominantly covered in the ground truth image. In an imbalanced feature distribution, the training process is therefore biased towards the larger feature subsets and colors of smaller objects evaporate from the resulting models. Therefore, the distribution of a*b* values is skewed towards desaturated values, and the color of minuscule objects disappears.
@@ III-C1 Continuous Color Range to Discrete Color Classes
The a* and b* color channels are continuous within the range $[-128,127]$. Each a*b* pair with a lightness value $L$ forms a RGB color pixel. We can get a a*b* pair from a*b* color space that is a 2-D space, where a* is one direction and b* is another direction. For a fixed $L$, small change in a*b* pair has no psychovisual effect. Therefore, human perception of the information in an image normally does not involve quantitative analysis of every pixel value in the image. As we know colorization is a regression problem where the regression model predicts the continuous quantities of a* and b* for a given $L$. Taking the advantage of psychovisual nature of human, the colorization problem can be represented as a classification problem where the learning model predicts a discrete class level for a a*b* pair. To fomularize the problem, the a*b* color space is divided into bins of a fixed grid size and each bin is assigned a discrete class level. The formula is given below in Equation 9.
@@
where a* and b* are the continuous color channels, $C$ is the discrete class level of color values within a bin, $\alpha^{2}$ is the area of a bin, $\beta$ is a shifting constant that shift a*b* color values into positive quadrant, $\delta$ is the number of grids in each a* or b* color channel.
bin size = 14 Refer to caption bin size = 12 Refer to caption bin size = 10 Refer to caption bin size = 8 Refer to caption bin size = 6 Refer to caption bin size = 4
@@ III-C2 Color Class to Visual Color Mapping
We need to extract a*b* pairs from the predicted color classes, $C$s, generated by the learning model for color image generation. Each bin is assigned by a fixed color class level $C$ which is driven by $a^{*}$ and $b^{*}$. The formulas are given below in Eq. 10, 11, which are the reverse of Eq. 9.
@@
According to the above equations, the maximum loss for each a* or b* value is $\frac{\alpha}{2}-1$. The higher value of $\alpha$ reduces the number of classes but makes the representation lossy as a large continuous range is converted to a single class. However, handling the problem with lower class is easy. The lower value of $\alpha$ increases the number of classes. In general, there is no guarantee that increasing or reducing the number of classes will increase classification accuracy. In colorization problem, more classes make the prediction less precise. It is important to adjust the number of class levels for a*b* color space so that modified $a^{*}b^{*}$ can describe image’s color nature.
@@ III-C3 Color Class Reduction Based on Practical Appearance
The a* and b* color values are continuous within $[-128,127]$ in the a*b* color channel. However, in practice, the range is found within $[-100,100]$. We first transform the continuous $[-108,108)$ ranged a*b* color channel to a single plane of 1296 color classes by taking $\alpha=6$, $\beta=108$ and $\delta=36$ in Eq. 9. A 2D grid of bins as $36\times 36$ single plane array is then formed where horizontal axes indicate a* and vertical axes indicate b* color information. Each coordinate is assigned a class value. The class matrix is shown in Figure 8. The proposed color class to visual color is shown in Figure 9. From Figure 9 we see that the whole image has a smooth representation of different colors. Colors of the nearest bins or blocks are almost similar. Color changes gradually block by block. We choose bin size 6 as we find it optimal for the colorization task. High bin size generates less class points and requires less model parameter. However, it increases the maximum and average deviation from the actual pixel during color values retrieval. Low bin size decreases the deviation but generates much classes and requires much model parameter. The bin size approximation is shown in Tab. II, and relative comparison is shown in Figure 13.
@@
With 1296 color classes we experimented with the Place365 Validation dataset [32] as a reference dataset. We see that some color classes face starvation in real-time images. The result is shown in Figure 10. In the Place365 Validation dataset, there are 36,500 images in total. We took 35040 images randomly from the dataset and extracted classes from them. We downsampled the images to $56\times 56$ to reduce the number of class samples in the training process. Therefore, there were a total of 109885440 (109.88M) class samples with 1296 class levels. Each color pixel is a color class sample with a specific color level. We calculate class level histogram on total class samples. We consider a class level for training samples if the class level has minimum of 500 (0.000455%) class samples out of 109885440 class samples. The class samples less than 500 with a specific class level is mapped to their nearest-neighbor present class levels using fixed centroid $k$-means clustering which is shown in Eq. 12. The class in the final color bin, which is shown in Figure 11. 500 pixels of a color is very minimum compared to the total pixels of 35040 images because very few images have very few number of that color and the most of the images do not contain this color. Finally we get 532 color classes within 1296 color classes which have more than 500 pixels. Class optimization is a major issue for colorization model. Less class may make the model more error-free, but it makes some color visuals outside of the bin. As one of our targets is to keep the rare appeared color values in the predicted distribution, we need to keep those rare color visuals active in the training process.
@@
where $C$ is the input color class vector, $\mu$ is the approved color classes for training, and $k$ is the number of color classes (532). We can define $\mu$ as the fixed value centroid. The iteration will happen single time and the centroid value will be unchanged.
@@ III-C4 Network Architecture
We build our model based on an encoder-decoder architecture. For the encoder part, we use DenseNet for our feature extractor. The DenseNet is a high-level feature extractor that is suitable for good color value generation. For the decoder part, we use conventional CNN. The network architecture of our proposed method is shown in Figure 12.
Figure 12: Network architecture of CCC++
@@
  • Colorization Network The network employs several convolutional and up-sampling layers after receiving an input of a $\frac{H}{32}\times\frac{W}{32}\times 1024$ feature representation. The fundamental nearest-neighbor method is what we employ for up-sampling. The $H\times W\times 2$ a*b* tensor is what the network outputs. In Table IV, the various convolutional layers and their results are displayed.
@@
III-C5 Loss Calculation
Colorization is generally considered a regression problem as the color values are continuous. But we transform the continuous color values into the discrete color classes described in Section III-C1. Therefore, we consider the problem as a classification problem and use cross-entropy loss instead of MSE or other regression loss. The loss function is shown in Equation 13.
Where $H$ and $W$ are the height and width of output $\kappa$ distribution. $\kappa$ is the true color class and $\overline{\kappa}$ is the estimated color class. The $\omega_{c}$ is the weights vector of color classes. The $\omega_{c}$ is defined as follows in Equation 14.
where $N_{c}$ is the number of color classes.
@@ III-C6 Class Confusion Based Weights Adjustment
In realistic images, all color classes are not represented equally. Grayish visual color classes are found in a much larger proportion than bright color classes due to the large background areas. In the categorical cross-entropy loss, each true class gets $\frac{1}{N}$ weight during loss calculation, which is shown in Eq. 14. As the bright color classes are far smaller in the count values, the gradients disappear gradually iteration by iteration. To keep the rarely appearing color classes, we increase the weights of the rarely appearing color classes more than the mostly appearing color classes. However, this process increases the global loss. So the weights need to be trade-offs to ensure both plausible colors and a minimum loss. To trade off the weights, we proposed a new formula, which is given in Equation 15.
@@
where $\mathcal{H}$, $\mathcal{W}$, $\mathcal{B}$ denote the height, width and batch size of the true class respectively. $\mathcal{C}$ is the color class set, ${\max}(N_{c})$ is the maximum appearance value of a class, $N_{c}$ is the appearance value of class $c$, $\mathcal{W}_{new}$ is the new weights matrix of the particular batch, $N^{adj}_{c}$ is the adjusted appearance value of class $c$, $\Psi$ is the threshold value for adjusting the value of appearance of the classes, $n_{c}$ is the total class number (532 is set by rule) in the optimized class set, $\mathcal{P}$ is the percentage value for determining the $\Psi$.
Initially, we regularize weights by dividing the total count of each class in a batch by its individual appearance value. This ensures that the weight of the most frequently appeared class is set to a minimum while proportionally upscaling the weights of others. However, this approach leads to a significant increase in the weight of classes appearing very infrequently. When a particular class appear 0 or tends to 0 then the weight of that class will be extremely high that makes the learning process imbalance again. To strike a balance, we introduce a trade-off mechanism. Firstly, we adjust appearance value of class $c$ using a threshold factor $\Psi$. By this we prevent every class for being holding extreme high weight. However, some minor classes appear very few times compared to some major classes. Therefore, those major classes go under the domination of some minor classes. To address the problem, we propose an additional term ($\frac{\max(N_{c})}{\Psi}$) for adding to the $N^{adj}_{c}$ term. When the difference between the maximum appeared class and minimum appeared $(\Psi)$ class is very high, the ratio will also be high. Therefore, high value will be added to the adjusted class value. This reduce the weights high appeared classes little but low appeared classes much. This seems little contradictory with our initial target as we are trying to establish the presence of minor class in the generated image and therefore, increasing the weights of the minor classes. However, unrestrained increase of weights of minor class can bias the learning towards minor classes. Our target is to ensure the learning unbiased. Therefore we need a perfect trade-off between the major and minor classes. To address this, we introduce a new hyperparameter $\mathcal{P}$ which denotes how much percentage we will take for determining the value of minimum appeared class. This supplementary term helps control the influence of rare class occurrences, providing a more nuanced and balanced approach to class weight determination. Therefore, the loss function is now modified, as shown in Equation. 16.
$Loss_{\text{CBCE}}=-\sum_{H,W}\mathcal{W}_{new}\sum_{C}\mathcal{K}.log(\overline{\mathcal{K}}).$ (16)
@@ III-C8 Class Selection from Probability Distribution for Image Reconstruction
Here, $\quad\mathcal{C}_{ij}$ represents the element at the i-th row and j-th column of matrix $\mathcal{C}$, and it is the index of the maximum probvability along the third dimension of matrix $P$ at the corresponding position.
     Gray      bin size = 4    bin size = 6    bin size = 8    bin size = 10    bin size = 12    bin size = 14    Ground Truth
@@ III-D SAM Empowered Object-Selective Color Harmonization
As we force regularization of the minor class, there is sometimes a little noise at the object’s edge. To make the edge more polished, we proposed SAM-empowered object-selective color harmonization which is shown in Figure 16. The SAM is a segmentation model with zero-shot generalization to unfamiliar objects and images without additional training[53]. The SAM is capable of extracting precise masks for multiple objects from an image[54]. Our proposed algorithm is shown in Algorithm 1.

### v2

@@ III-A Color Space
Left: bin size = 10; Right: bin size = 8 Refer to caption Refer to caption
Left: bin size = 6; Right: bin size = 4
@@ III-B Problem Definition
The colorization problem involves predicting color channels based on a given gray channel. The Lightness (L) channel of the La*b* color space can be converted to the gray channel (intensity) and vice versa[61]. Additionally, it is possible to convert RGB values into LAB color space and vice versa. The task can be defined as follows in Equation 1, 2, 3, and 4.
@@
where $X_{L}$ represents the lightness channel, $X_{ab}$ represents the predicted color channel, $X_{Lab}$ represents the predicted color image, $Y_{ab}$ represents the ground truth color channel. The function $f(.)$ represents the mapping function achieved through deep learning. The function $Distance_{min}(.)$ represents the objective function, which can be any loss function the optimizer uses to make the learning process efficient. The set $\mathbb{R}$ represents the entirety of the image component, whereas $\mathbb{H}$ and $\mathbb{W}$ represent the dimensions of the image.
@@
Theoretically, the values of the a* and b* channels range from -128 to 127, inclusively, and are continuous. Therefore, the task at hand is classified as a regression problem. Hence, the $Distance_{min}(.)$ in equation 2 can be expressed as either L1 loss, L2 loss, Huber loss, Log-cosh loss, or a similar regression loss. The loss functions are displayed in Equation 5, 6, 7 and , 8.
@@
The variable $N$ represents the total number of pixels in a batch, whereas the variable $\delta$ represents the acceptable threshold. The real-time photographs predominantly feature background colors that resemble clouds, earth, pavement, and walls. The color components of the major or focal objects are significantly less than the total color components, which leads to an uneven allocation of characteristics. Addressing feature imbalance is essential as the smaller groups of features are the focal point for the learning process. The intrinsic ambiguity and multimodality of the colorization problem render the above loss functions susceptible. If an object can have different values for a*b*, the most efficient way to solve the loss indicated in equations 5, 6, 7, and 8 is by calculating the mean of the set. The averaging error effect prefers color values that are mostly present in the ground truth image. When there is an uneven distribution of features, the training process becomes biased towards the larger sets of features. As a result, the models generated from this method lose the colors of smaller objects. Hence, the distribution of a*b* values is biased towards desaturated values, resulting in the disappearance of color in tiny objects.
@@ III-C Solution Approach
The system architecture of our proposed approach is shown in Fig. 7. The details of our proposed model is described below.
Left: bin size = 10; Right: bin size = 8 Refer to caption Refer to caption
Left: bin size = 6; Right: bin size = 4
@@ III-C1 Continuous Color Range to Discrete Color Classes
The color channels a* and b* exhibit continuous values throughout the range of -128 to 127. An RGB color pixel is formed by each pair of a*b* values with a lightness value $L$. A pair of a*b* can be obtained from a*b* color space, a two-dimensional space where a* represents one direction and b* represents another. For a constant value of $L$, a slight alteration in the pair of variables a and b produces no noticeable visual impact. Thus, the human perception of information in a picture typically does not entail a quantitative examination of each pixel value inside the image. Colorization is a regression problem in which a regression model predicts the continuous values of a* and b* based on a given $L$. By exploiting the psycho-visual characteristics of humans, the colorization problem can be framed as a classification problem in which the learning model predicts a discrete class level for a given a*b* pair. The a*b* color space is partitioned into bins of predetermined grid size, and each bin is assigned a discrete class level to formulate the problem. The formula is provided in Equation 9.
@@
where a* and b* represent the continuous color channels, $C$ represents the discrete class level of color values within a bin, $\alpha^{2}$ represents the area of a bin, $\beta$ represents a shifting constant that shifts a*b* color values into a positive quadrant, $\delta$ represents the number of grids in each a* or b* color channel.
@@ III-C2 Color Class to Visual Color Mapping
We need to retrieve pairs of a*b* values from the color classes, $C$s, predicted by the color generation learning model. Each bin is allocated a specific color class level $C$ based on the values of $a^{*}$ and $b^{*}$. The formulas are provided below as Equation 10 and 11, which represent the inverse of Equation 9.
@@
Based on the given equations, the maximum loss for each value of a* or b* equals $\frac{\alpha}{2}-1$. Increasing the amount of $\alpha$ decreases the number of classes, but it also causes the representation to become lossy as a broad continuous range is compressed into a single class. Nevertheless, addressing the issue concerning the lower socioeconomic class is a straightforward task. Decreasing the value of $\alpha$ leads to an increase in the number of classes. Generally, there is no assurance adjusting the number of classes would enhance or diminish classification accuracy. Increasing the number of classes in the colorization problem leads to decreased prediction accuracy. It is crucial to calibrate the number of class levels for the a*b* color space to accurately represent the image’s color characteristics using the updated $a^{*}b^{*}$ values.
@@ III-C3 Color Class Reduction Based on Practical Appearance
The color values a* and b* range continuously from -128 to 127 in the a*b* color channel. Nevertheless, in practical terms, the range is generally confined to between -100 to 100. The continuous color channel a*b* in the range of $[-108,108)$ is converted into a single plane consisting of 1296 color classes. This transformation is achieved by using the values $\alpha=6$, $\beta=108$, and $\delta=36$ in Equation 9. A 2D grid of bins is created by arranging a single plane array of size $36\times 36$. The horizontal axes represent the* color information, while the vertical axes represents the b* color information. A class value is assigned to each coordinate. The matrix representing the class is displayed in Fig. 8. The color class proposed for visualizing color is depicted in Fig. 9. Fig. 9 demonstrates that the entire image seamlessly depicts various hues. The colors of the neighboring bins or blocks are nearly the same. The color transitions occur progressively, one block at a time. We choose bin size 6 as we find it optimal for the colorization task. High bin size generates less class points and requires less model parameter. However, it increases the maximum and average deviation from the actual pixel during color values retrieval. Low bin size decreases the deviation but generates much classes and requires much model parameter. The bin size approximation is shown in TABLE II, and relative comparison is shown in Fig. 13.
Figure 12: Network architecture of CCC++
@@
With 1296 color classes we experimented with the Place365 Validation dataset [55] as a reference dataset. We see that some color classes face starvation in real-time images. The result is shown in Fig. 10. In the Place365 Validation dataset, there are 36,500 images in total. We randomly selected 35040 photos from the dataset and extracted classes from each of them. To decrease the number of class samples during the training phase, we resized the images to dimensions of $56\times 56$. Consequently, there was a total of 109,885,440 (109.88 million) class samples, each belonging to one of the 1,296 class levels. Every pixel represents a color class sample with a distinct color intensity. We calculate class level histogram on total class samples. We consider a class level for training samples if the class level has minimum of 500 (0.000455%) class samples out of 109885440 class samples. The class samples with a class level below 500 are assigned to their nearest-neighbor class levels using fixed centroid $k$-means clustering, as described in Equation 12. The class in the final color bin, which is shown in Fig. 11. 500 pixels of a color is very minimum compared to the total pixels of 35040 images because very few images have very few number of that color and the most of the images do not contain this color. In the end, we obtained 532 color classes out of a total of 1296 color classes that contain more than 500 pixels. Efficient optimization of the class is a significant concern for the colorization model. Reducing the number of classes in the model may improve its accuracy by reducing errors, but it may also result in some color visuals appearing outside of the designated range. One of our objectives is to maintain the presence of infrequently occurring color values in the projected distribution. Therefore, it is necessary to ensure that these rare color visions remain active during the training process.
@@
$C$ represents the input color class vector, $\mu$ represents the approved color classes for training, and $k$ represents the number of color classes equal to 532. The term $\mu$ can be defined as the stationary centroid value. The loop will occur only once, and the centroid value will remain unchanged.
@@ III-C4 Network Architecture
Our model is constructed using an encoder-decoder architecture. Our feature extractor for the encoder part is DenseNet. The DenseNet is a sophisticated feature extractor that is well-suited for generating high-quality color values. Conventional CNN is employed for the decoder component. The network structure of our suggested approach is illustrated in Fig. 12.
@@
  • Colorization Network After obtaining a feature representation of size $\frac{H}{32}\times\frac{W}{32}\times 1024$, the network leverages many convolutional and up-sampling layers. We utilize the fundamental nearest-neighbor method for up-sampling. The network outputs a tensor of dimensions $H\times W\times 2$ with elements denoted by a*b*. TABLE IV presents the convolutional layers and their corresponding outcomes.
@@
III-C5 Loss Calculation
Colorization is often regarded as a regression problem due to the continuous nature of color values. However, we convert the continuous color values into the discrete color classes outlined in Section III-C1. Thus, we approach the problem as a classification problem and employ cross-entropy loss instead of mean squared error (MSE) or other regression loss functions. The loss function is depicted in Equation 13.
Here, $H$ and $W$ represent the height and width of the output $\kappa$ distribution. $\kappa$ represents the actual color category, while $\overline{\kappa}$ represents the predicted color category. The $\omega_{c}$ represents the vector of weights for color classes. The symbol $\omega_{c}$ is defined in Equation 14.
where $N_{c}$ is the number of color classes.
@@ III-C6 Class Confusion Based Weights Adjustment
In actual photographs, the distribution of color classes is not uniform. The prevalence of grayish visual color classes is higher than luminous color classes, primarily because of the extensive background areas. The categorical cross-entropy loss assigns a weight of $\frac{1}{N}$ to each actual class when calculating the loss, as seen in Equation 14. Due to the relatively low frequency of bright color classes, the gradients steadily diminish with each iteration. To preserve the infrequently occurring color categories, we amplify the weights of these categories to a greater extent than those of the predominantly occurring color categories. Nevertheless, this procedure amplifies the overall detriment on a worldwide scale. Making trade-offs while determining the weights is necessary to achieve both realistic colors and minimize loss. We have introduced a novel formula to balance the weights, as shown in Equation 15.
@@
where $\mathcal{H}$, $\mathcal{W}$, and $\mathcal{B}$ represent the height, width, and batch size of the true class, respectively. $\mathcal{C}$ represents the set of color classes. ${\max}(N_{c})$ refers to a class’s highest value of appearance. $N_{c}$ represents the appearance value of a specific class $c$. $\mathcal{W}_{new}$ is the matrix of weights for the current batch. $N^{adj}_{c}$ represents the adjusted appearance value of a class. The variables $c$ and $\Psi$ represent the threshold value used to change the appearance value of the classes. The variable $n_{c}$ represents the total number of classes in the optimized class set, which is set to 532 according to a rule. The variable $\mathcal{P}$ represents the percentage value used to determine the threshold value $\Psi$.
@@ G. Truth
Initially, we regularize weights by dividing the total count of each class in a batch by its individual appearance value. This ensures that the weight of the most frequently appeared class is set to a minimum while proportionally upscaling the weights of others. However, this approach leads to a significant increase in the weight of classes appearing very infrequently. When a particular class appear 0 or tends to 0 then the weight of that class will be extremely high that makes the learning process imbalance again. To strike a balance, we introduce a trade-off mechanism. Firstly, we adjust appearance value($N^{adj}_{c}$) of class $c$ using a threshold factor $\Psi$. By this we prevent every class for being holding extreme high weight. However, some minor classes appear very few times compared to some major classes. Therefore, those major classes go under the domination of some minor classes. To address the problem, we propose an additional term ($\frac{\max(N_{c})}{\Psi}$), to be incorporated into the $N^{adj}_{c}$ term. When the difference between the maximum appeared class and minimum appeared $(\Psi)$ class is very high, the ratio will also be high. Therefore, high value will be added to the adjusted class value. This reduce the weights high appeared classes little but low appeared classes much. This seems little contradictory with our initial target as we are trying to establish the presence of minor class in the generated image and therefore, increasing the weights of the minor classes. However, unrestrained increase of weights of minor class can bias the learning towards minor classes. Our target is to ensure the learning unbiased. Therefore we need a perfect trade-off between the major and minor classes. To address this, we introduce a new hyperparameter $\mathcal{P}$ which denotes how much percentage we will take for determining the value of minimum appeared class. This additional term assists in managing the impact of infrequent class instances, resulting in a more refined and equitable method for determining class weights. Consequently, the loss function has been altered, as depicted in the Equation 16.
$\vspace{-4mm}Loss_{\text{CBCE}}=-\sum_{H,W}\mathcal{W}_{new}\sum_{C}\mathcal{K}.log(\overline{\mathcal{K}}).\vspace{-1mm}$ (16)
@@ III-C8 Class Selection from Probability Distribution for Image Reconstruction
Here, $\quad\mathcal{C}_{ij}$ represents the element at the i-th row and j-th column of matrix $\mathcal{C}$, and it is the index of the maximum probability along the third dimension of matrix $P$ at the corresponding position.
Gray    bs=4   bs=6   bs=8   bs=10   bs=12   bs=14   G. Tru.
@@ III-D SAM Empowered Object-Selective Color Harmonization
To address the noise sometimes observed at the edges of objects when regularizing the minor class, we propose SAM-empowered object-selective color harmonization, as illustrated in Fig. 16. The SAM is a cutting-edge image segmentation model developed by Meta AI Research. It features zero-shot generalization, allowing it to segment various objects and scenes without additional training. SAM excels in interactive segmentation, enabling users to refine results with points, boxes, or text prompts. Known for its high precision, SAM can extract detailed masks for multiple objects in complex images, making it highly versatile and effective for various applications [63, 64]. Our proposed algorithm is detailed in Algorithm 1. This approach ensures more polished edges by leveraging SAM’s advanced segmentation capabilities, leading to improved overall color harmonization.